# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ 2 Computing spectral projectors with the sign function
The full proof of Proposition 2.1 can be found in Appendix B. We briefly summarize it here. The main idea is to use the sign function algorithm from [BGVKS22], $\mathsf{SGN}$, to approximate $\operatorname{sgn}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. If we already know that $\widetilde{\mu}_{k}$ is a reasonable approximation of $\mu_{k}$, and that it is located inside the correct eigenvalue gap, then, in exact arithmetic, our problem is equivalent to computing $\operatorname{sgn}(\mu_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. The result can be used to filter the desired spectral projector, often referred as “purification” in the context of DFT. The main challenge is to ensure that all propagated numerical errors, success probabilities, and input parameters for all algorithms are well-defined and bounded. To obtain the final forward errors we must rely on matrix similarity arguments, the properties of the pseudospectrum, the eigenvalue bounds of Weyl and Kahan from Fact A.1, and Lemma B.1, which gives explicit bounds on the sign function under small floating point perturbations. ∎
The rest of the paper is devoted to the analysis of our new algorithm for the spectral gap and the midpoint based on eigenvalue counting queries, described in Theorem 3.1. For comparison purposes, in Appendix B.1 we analyze a diagonalization-based algorithm for the same task (which is a new result itself), specifically, using the state-of-the-art $\mathsf{EIG}$ algorithm of [BGVKS22]. We compare the two algorithms in Section 3.4, demonstrating that our counting-based algorithm is indeed faster.
@@ 3.1 Smoothed analysis of eigenvalue counting
The main result that we use in the proof can be traced back to the Wegner estimate for the density-of-states of Hermitian operators under random diagonal disorder [Weg81]. See Appendix D and in particular D.2 for more details. ∎
@@ 3.2 Computing the gap and the midpoint
Let $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$ be a definite pencil with $\|\boldsymbol{\mathrm{H}}\|,\|\boldsymbol{\mathrm{S}}^{-1}\|\leq 1$. Given $k\in[n-1]$, accuracy $\epsilon\in(0,1)$, and failure probability $\delta\in(0,1/2)$, there exists an algorithm
@@
which returns $\widetilde{\mu}_{k}=\mu_{k}\pm\epsilon\operatorname{gap}_{k}$ and $\widetilde{\operatorname{gap}}_{k}=(1\pm\epsilon)\operatorname{gap}_{k}$, where $\mu_{k}=\tfrac{\lambda_{k}+\lambda_{k+1}}{2}$, $\operatorname{gap}_{k}=\lambda_{k+1}-\lambda_{k}$, and $\lambda_{i}$ are the eigenvalues of $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$. The algorithm requires $O\left(T_{\mathsf{MM}}(n)\log(\tfrac{1}{\delta\epsilon\operatorname{gap}_{k}})\log(\tfrac{1}{\epsilon\operatorname{gap}_{k}})\right)$ arithmetic operations using $O\left(\log(n)\log(\tfrac{1}{\delta\epsilon\operatorname{gap}_{k}})\log^{3}(\tfrac{n}{\delta\epsilon\operatorname{gap}_{k}})\right)$ bits.
@@ 3.4 Comparison with diagonalization
We can now compare Theorem 3.1 with a diagonalization-based approach that is detailed in Proposition B.2. For $\epsilon,\operatorname{gap}_{k},\kappa^{-1}(\boldsymbol{\mathrm{S}})\in\Omega(\operatorname{poly}(1/n))$, the total arithmetic complexity of the algorithm of Theorem 3.1 is $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ using $O(\log^{5}(n))$ bits. For the same parameters, Proposition B.2 requires need a total of $O(T_{\mathsf{MM}}(n)\log^{3}(n))$ arithmetic operations, and $O(\log^{5}(n))$ bits. Thus, in total, Algorithm 3.1 is faster by a factor of $O(\log(n))$.
In the extreme case where $\epsilon,\operatorname{gap}_{k},\kappa(\boldsymbol{\mathrm{S}})=\Theta(1)$, Theorem 3.1 counts $O(T_{\mathsf{MM}}(n)\log(n))$ arithmetic operations and $O(\log^{5}(n))$ bits, while Proposition B.2 requires $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ operations, and $O(\log^{5}(n))$ bits. Thus, Proposition B.2 is again slower by a factor of $O(\log(n))$. Interestingly, in this case Theorem 3.1 is faster than even a single call to $\mathsf{EIG}$, which requires $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ arithmetic operations. We conclude that, at least based on the currently existing algorithms, diagonalization is redundant for the computation of invariant subspaces.
@@ 3.5 Application in DFT
In Appendix F we demonstrate how our main results can be directly applied to approximate density matrices and electron densities of atomic systems in DFT. Even though is not a machine learning problem per se, we decided to dedicate a section in the Appendix due to its importance: DFT calculations persistently occupy supercomputing clusters and the corresponding software libraries and literature receive tens of thousands of citations annually at the time of this writing [KF96, GAB+17, SAG+02, HISV14]. Our work is the first analysis to provide forward-error guarantees in finite precision for these problems in nearly matrix multiplication time.
@@ 4 PCA
Since its introduction in the early twentieth century [Pea01, Hot33], Principal Component Analysis is one of the most important tools in statistics, data science, and machine learning. It can be used, for example, to visualize data, to reduce dimensionality, or to remove noise from data; cf. [Jol02, DK96] for reviews on the vast bibliography. In its simplest formulation, given a (centered) data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, the goal is to learn a $k$-dimensional embedding $\boldsymbol{\mathrm{C}}_{k}$, where $k<n$, that maximizes the sample variance, which can be written as an optimization problem
@@
(which is indeed a definite GEP as in Equation (1) with $\boldsymbol{\mathrm{S}}=\boldsymbol{\mathrm{I}}$). By solving for $\boldsymbol{\mathrm{C}}_{k}$, we can project the data in $k$ dimensions by computing $\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{C}}_{k}$, preserving as much of the variance in $k$ dimensions as possible. To compute $\boldsymbol{\mathrm{C}}_{k}$ we can directly use our main results. However, the solution of Equation (3) is an actual orthonormal basis for the invariant subspace rather than the spectral projector that Theorem 1.1 returns. This can be addressed with deflation: Once we have the spectral projector $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$, assuming that the approximation is sufficiently tight, we can apply a subsequent deflation step based on a rank-revealing QR factorization to obtain a $k$-dimensional basis. This can be done deterministically in $O(n^{3})$ time [GE96] or in randomized $O(n^{\omega})$ [DDH07, BGVKS22].
In Appendix G.1 we prove the following Theorem 4.1 for Algorithm 7, which builds upon our main Theorem 1.1, the algorithm of Proposition E.2 to approximate $\|\boldsymbol{\mathrm{X}}-\boldsymbol{\mathrm{X}}_{k}\|$, and the $\mathsf{DEFLATE}$ algorithm of [BKMS21], to solve the standard PCA problem of Eq. (3). Following the existing literature, the result is stated for real matrices, but it can be trivially adapted to the complex case as well.
@@ 4.1 Block-Krylov PCA
In some applications, the target dimension $k$ might be small, i.e., $k\ll n$. This condition has driven a whole area of research in so-called low-rank approximation algorithms for PCA [FKV98, Sar06, CW09, HMT11, MRT11, MM15, BDMI14, BJS14, AZL16]. Such approaches are also suitable for kernel PCA, since they rely on matrix-vector products and therefore the kernel matrix does not need to be explicitly formed. The techniques from the previous section can be directly applied to obtain new bit complexity upper bounds for existing state-of-the-art algorithms, which are typically analyzed in exact arithmetic. They internally rely on the computation of the principal singular vectors of submatrices, which can be improved with our methods. Specifically, in Appendix G.2 we summarize a floating point analysis of the Block Krylov Iteration algorithm (see Algorithm 8), essentially, providing a matrix multiplication-type upper bound on the bit complexity with only a polylogarithmic dependence on the singular value gap. In a nutshell, we directly obtain the following corollary:
Corollary 4.1 (Bit complexity analysis of Block Krylov PCA).
Let $\boldsymbol{\mathrm{X}}$ be a data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, $\|\boldsymbol{\mathrm{X}}\|\leq 1$, $k\in[n]$ is a target rank, and $\epsilon_{\mathsf{PCA}}\in(0,1)$ an accuracy parameter. Let $q=\frac{\sqrt{\log(n)}}{\epsilon_{\mathsf{PCA}}}$. Using the Steps 1-6 that are detailed in Appendix G.3 as a floating point implementation of Algorithm 8, we can compute with high probability a matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\in\mathbb{R}^{m\times k}$ that satisfies
$\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\|\leq O(\epsilon_{\mathsf{PCA}}),$
where $\boldsymbol{\mathrm{Z}}_{k}$ is an approximate basis for the top-$k$ principal components of $\boldsymbol{\mathrm{X}}$, returned by Algorithm 8 in exact arithmetic. The total cost is at most
$O\left(qmnk^{\omega-2}\operatorname{polylog}(\tfrac{qmnk\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mnqk\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision.
@@
The full proof can be found in Appendix G.3. The main idea is to apply the counting query methodology to compute the condition number of the Block-Krylov matrix, as well as the $k$-th spectral gap and the midpoint of the reduced matrix $\boldsymbol{\mathrm{M}}$ in Line 5 of Algorithm 8. Thereafter, we can compute a spectral projector and an approximate basis for the top-$k$ singular vectors of $\boldsymbol{\mathrm{M}}$ using $\mathsf{PURIFY}$ and $\mathsf{DEFLATE}$, similar to the analysis of classical PCA in the previous section. ∎
@@ 5 Open problems
    Bit requirement of $\mathsf{SGN}$: The major bottleneck for the bit requirements of our main algorithms comes from the $\mathsf{SGN}$ algorithm of [BGVKS22]. An inverse-free Newton-Schultz iteration [KL95], or the implicit repeated squaring of [DDS23, Sch23] can potentially give significant improvements.
@@
    Fast rectangular matrix multiplication: In some parts of the analysis we used blocked-square matrix multiplication to multiply rectangular matrices. Taking advantage of rectangular matrix multiplication [LG12, Gal24] can potentially lead to faster algorithms.
@@
    Sparse algorithms: In applications like DFT it commonly occurs that the matrices have special structure, i.e., they are banded and/or sparse. It remains open whether Problem 1.1 can be provably solved faster than our reported results in finite precision for these special cases (recall that the tridiagonal QR algorithm requires $O(n^{3})$ operations to return the eigenvectors). An end-to-end stability analysis of existing fast eigensolvers would be the place to start [GE95, VXCB16].
  4. (iv)
    Distributed PCA: The techniques for Block-Krylov PCA can be potentially applied to distributed or streaming PCA algorithms, which are also based on randomized low-rank approximations. E.g., in the distributed PCA algorithm of [BWZ16], it is straightforward to replace the SVD computation on the server by a counting query iteration. The full analysis of such an approach is left as future work.
@@ Appendix A Preliminaries
In this section the model of computation is defined and basic linear algebra principles are summarized. Further details can be found in standard textbooks such as [Hig02, GVL13].
@@ A.1 Forward and backward approximation
In numerical analysis, a “forward-error” often measures the distance between the true solution to a problem and the solution returned by a numerical algorithm. A related notion is the so-called “backward error” or “backward approximation.” For more details we can refer to the standard textbook of Higham [Hig02]. In this case, the solution returned is the exact solution of a “nearby problem,” and the backward error quantifies the distance of the original problem to this nearby problem. In PCA, for example, a backward-approximation could be defined by $\widetilde{\boldsymbol{\mathrm{C}}}_{k}=\arg\min_{\boldsymbol{\mathrm{C}}^{\top}\boldsymbol{\mathrm{C}}=\boldsymbol{\mathrm{I}}_{k\times k}}\operatorname{tr}(\boldsymbol{\mathrm{C}}^{\top}\widetilde{\boldsymbol{\mathrm{H}}}\boldsymbol{\mathrm{C}})$. The backward error would be given $\|\boldsymbol{\mathrm{H}}-\widetilde{\boldsymbol{\mathrm{H}}}\|$ for some norm. An algorithm is backward stable if the backward error is always well-defined and bounded. A forward error type of approximation is often harder to achieve. A common rule-of-thumb states that
@@
However, this does not generally hold for eigenvalue problems. For example, Proposition 1.1 in [BGVKS22] can be used to transformed a backward approximate diagonalization error to a forward error for the eigenvectors, but the bound depends on the minimum eigenvalue gap between any eigenvalue pair. As the authors point in Remark 1.2, special treatment is needed in terms of the invariant subspaces in the presence of multiple eigenvalues, which are thoroughly analyzed in this work.

### v4

@@ 2 Computing spectral projectors with the sign function
The full proof of Proposition 2.1 can be found in Appendix B. We briefly summarize it here. The main idea is to use the sign function algorithm from [15], $\mathsf{SGN}$, to approximate $\operatorname{sgn}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. If we already know that $\widetilde{\mu}_{k}$ is a reasonable approximation of $\mu_{k}$, and that it is located inside the correct eigenvalue gap, then, in exact arithmetic, our problem is equivalent to computing $\operatorname{sgn}(\mu_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}})$. The result can be used to filter the desired spectral projector, often referred as “purification” in the context of DFT. The main challenge is to ensure that all propagated numerical errors, success probabilities, and input parameters for all algorithms are well-defined and bounded. To obtain the final forward errors we must rely on matrix similarity arguments, the properties of the pseudospectrum, the eigenvalue bounds of Weyl and Kahan from Fact A.1, and Lemma B.1, which gives explicit bounds on the sign function under small floating point perturbations. ∎
The rest of the paper is devoted to the analysis of our new algorithm for the spectral gap and the midpoint based on eigenvalue counting queries, described in Theorem 3.1. For comparison purposes, in Appendix B.1 we analyze a diagonalization-based algorithm for the same task (which is a new result itself), specifically, using the state-of-the-art $\mathsf{EIG}$ algorithm of [15]. We compare the two algorithms in Section 3.4, demonstrating that our counting-based algorithm is indeed faster.
@@ 3.1 Smoothed analysis of eigenvalue counting
The main result that we use in the proof can be traced back to the Wegner estimate for the density-of-states of Hermitian operators under random diagonal disorder [138]. See Appendix D and in particular D.2 for more details. ∎
@@ 3.2 Computing the gap and the midpoint
arithmetic operations using $O\left(\log(n)\left(\log^{4}(\tfrac{n}{\delta\epsilon\operatorname{gap}_{k}})+\log(\kappa(\boldsymbol{\mathrm{S}}))\right)\right)$ bits, where $\lambda_{i}$ are the eigenvalues of $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$. If $\kappa(\boldsymbol{\mathrm{S}})$ is unknown, additional $O(T_{\mathsf{MM}}(n)\log(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta})\log(\kappa(\boldsymbol{\mathrm{S}})))$ floating point operations and $O(\log(n)\log^{4}(\tfrac{n\kappa(\boldsymbol{\mathrm{S}})}{\delta}))$ bits are sufficient to compute it with Corollary E.1.
@@ 3.4 Comparison with diagonalization
We can now compare Theorem 3.1 with a diagonalization-based approach that is detailed in Proposition B.2. We fix $\delta=O(1/n)$ so that both algorithms succeed with the same probability.
For $\epsilon,\operatorname{gap}_{k},\kappa^{-1}(\boldsymbol{\mathrm{S}})\in\Omega(\operatorname{poly}(1/n))$, the total arithmetic complexity of the algorithm of Theorem 3.1 is $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ using $O(\log^{5}(n))$ bits. For the same parameters, Proposition B.2 requires need a total of $O(T_{\mathsf{MM}}(n)\log^{3}(n))$ arithmetic operations, and $O(\log^{5}(n))$ bits. Thus, in total, Algorithm 3.1 is faster by a factor of $O(\log(n))$.
In the extreme case where $\epsilon,\operatorname{gap}_{k},\kappa(\boldsymbol{\mathrm{S}})=\Theta(1)$, Theorem 3.1 counts $O(T_{\mathsf{MM}}(n)\log(n))$ arithmetic operations and $O(\log^{5}(n))$ bits, while Proposition B.2 requires $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ operations, and $O(\log^{5}(n))$ bits. Thus, Proposition B.2 is again slower by a factor of $O(\log(n))$. Interestingly, in this case Theorem 3.1 is faster than even a single call to $\mathsf{EIG}$, which requires $O(T_{\mathsf{MM}}(n)\log^{2}(n))$ arithmetic operations. We conclude that, at least based on the currently existing algorithms, diagonalization is redundant for the computation of spectral gaps and invariant subspaces.
@@ 3.5 Application in DFT
In Appendix F we demonstrate how our main results can be directly applied to approximate density matrices and electron densities of atomic systems in DFT. Even though is not a machine learning problem per se, we decided to dedicate a section in the Appendix due to its importance: DFT calculations persistently occupy supercomputing clusters and the corresponding software libraries and literature receive tens of thousands of citations annually at the time of this writing [89, 59, 60, 127, 74]. Our work is the first analysis to provide forward-error guarantees in finite precision for these problems in nearly matrix multiplication time.
@@ 4 PCA
Since its introduction in the early twentieth century [116, 72], Principal Component Analysis is one of the most important tools in statistics, data science, and machine learning. It can be used, for example, to visualize data, to reduce dimensionality, or to remove noise from data; cf. [79, 45] for reviews on the vast bibliography. In its simplest formulation, given a (centered) data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, the goal is to learn a $k$-dimensional embedding $\boldsymbol{\mathrm{C}}_{k}$, where $k<n$, that maximizes the sample variance, which can be written as an optimization problem
@@
(which is indeed a definite GEP as in Equation (1) with $\boldsymbol{\mathrm{S}}=\boldsymbol{\mathrm{I}}$). By solving for $\boldsymbol{\mathrm{C}}_{k}$, we can project the data in $k$ dimensions by computing $\boldsymbol{\mathrm{X}}\boldsymbol{\mathrm{C}}_{k}$, preserving as much of the variance in $k$ dimensions as possible. To compute $\boldsymbol{\mathrm{C}}_{k}$ we can directly use our main results. However, the solution of Equation (3) is an actual orthonormal basis for the invariant subspace rather than the spectral projector that Theorem 1.1 returns. This can be addressed with deflation: Once we have the spectral projector $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$, assuming that the approximation is sufficiently tight, we can apply a subsequent deflation step based on a rank-revealing QR factorization to obtain a $k$-dimensional basis. This can be done deterministically in $O(n^{3})$ time [66] or in randomized $O(n^{\omega})$ [39, 15].
In Appendix G.1 we prove the following Theorem 4.1 for Algorithm 7, which builds upon our main Theorem 1.1, the algorithm of Proposition E.2 to approximate $\|\boldsymbol{\mathrm{X}}-\boldsymbol{\mathrm{X}}_{k}\|$, and the $\mathsf{DEFLATE}$ algorithm of [19], to solve the standard PCA problem of Eq. (3). Following the existing literature, the result is stated for real matrices, but it can be trivially adapted to the complex case as well.
@@ 4.1 Block-Krylov PCA
In some applications, the target dimension $k$ might be small, i.e., $k\ll n$. This condition has driven a whole area of research in so-called low-rank approximation algorithms for PCA [53, 120, 33, 67, 100, 106, 30, 29, 5]. Such approaches are also suitable for kernel PCA, since they rely on matrix-vector products and therefore the kernel matrix does not need to be explicitly formed. The techniques from the previous section can be directly applied to obtain new bit complexity upper bounds for existing state-of-the-art algorithms, which are typically analyzed in exact arithmetic. They internally rely on the computation of the principal singular vectors of submatrices, which can be improved with our methods. Specifically, in Appendix G.2 we summarize a floating point analysis of the Block-Krylov Iteration algorithm (see Algorithm 8), essentially, providing a matrix multiplication-type upper bound on the bit complexity with only a polylogarithmic dependence on the singular value gap. In a nutshell, we directly obtain the following result:
Theorem 4.2 (Bit complexity analysis of Block-Krylov PCA).
Let $\boldsymbol{\mathrm{X}}$ be a data matrix $\boldsymbol{\mathrm{X}}\in\mathbb{R}^{m\times n}$, $\|\boldsymbol{\mathrm{X}}\|\leq 1$, $k\in[n]$ a target rank, $\epsilon_{\mathsf{PCA}}\in(0,1)$ an accuracy parameter, and $q=\Theta\left(\frac{\log(n)}{\sqrt{\epsilon_{\mathsf{PCA}}}}\right)$. Let $T_{\mathsf{MMX}}(k)$ denote the complexity to stably multiply $\boldsymbol{\mathrm{X}}$ or $\boldsymbol{\mathrm{X}}^{\top}$ with a dense matrix with $k$ columns from the right (see Def. G.1). Using the Steps 1-6 that are detailed in Appendix G.3 as a floating point implementation of Algorithm 8, we can compute a matrix $\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\in\mathbb{R}^{m\times k}$ that satisfies
$\left\|\widetilde{\boldsymbol{\mathrm{Z}}}_{k}\widetilde{\boldsymbol{\mathrm{Z}}}_{k}^{\top}-\boldsymbol{\mathrm{Z}}_{k}\boldsymbol{\mathrm{Z}}_{k}^{\top}\right\|\leq O(\epsilon_{\mathsf{PCA}}),$
with high probability, where $\boldsymbol{\mathrm{Z}}_{k}$ is an approximate basis for the top-$k$ principal components of $\boldsymbol{\mathrm{X}}$, returned by Algorithm 8 in exact arithmetic. The total cost is at most
$O\left(qT_{\mathsf{MMX}}(k)\log(\tfrac{\kappa(\boldsymbol{\mathrm{K}})}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+m(qk)^{\omega-1}\log(\tfrac{1}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})+(qk)^{\omega}\operatorname{polylog}(\tfrac{qk}{\operatorname{gap}_{k}(\boldsymbol{\mathrm{M}})})\right)$
floating point operations, using $O\left(\operatorname{polylog}(\tfrac{mq\kappa(\boldsymbol{\mathrm{K}})}{\epsilon_{\mathsf{PCA}}\operatorname{gap}_{k}})\right)$ bits of precision. $\boldsymbol{\mathrm{K}},\boldsymbol{\mathrm{M}}$ are as in Alg. 8.
@@
The full proof can be found in Thm. G.2, Appendix G.3. The main idea is to apply the counting query methodology to compute the condition number of the Block-Krylov matrix $\boldsymbol{\mathrm{K}}$, as well as the $k$-th spectral gap and the midpoint of the reduced matrix $\boldsymbol{\mathrm{M}}$ in Line 5 of Alg. 8. Thereafter, we can compute a spectral projector and an approximate basis for the top-$k$ singular vectors of $\boldsymbol{\mathrm{M}}$ using $\mathsf{PURIFY}$ and $\mathsf{DEFLATE}$, similar to the analysis of classical PCA in the previous section. ∎
@@ 5 Open problems
    Bit requirement of $\mathsf{SGN}$: The major bottleneck for the bit requirements of our main algorithms comes from the $\mathsf{SGN}$ algorithm of [15]. An inverse-free Newton-Schultz iteration [83], or the implicit repeated squaring of [41, 121] can potentially give significant improvements.
@@
    Sparse algorithms: In applications like DFT it commonly occurs that the matrices have special structure, i.e., they are banded and/or sparse. It remains open whether Problem 1.1 can be provably solved faster than our reported results in finite precision for these special cases (recall that the tridiagonal QR algorithm requires $O(n^{3})$ operations to return the eigenvectors). An end-to-end stability analysis of existing fast eigensolvers would be the place to start [65, 137].
@@
    Distributed PCA: The techniques for Block-Krylov PCA can be potentially applied to distributed or streaming PCA algorithms, which are also based on randomized low-rank approximations. E.g., in the distributed PCA algorithm of [31], it is straightforward to replace the SVD computation on the server by a counting query iteration. The full analysis of such an approach is left as future work.
@@ Appendix A Preliminaries
In this section the model of computation is defined and basic linear algebra principles are summarized. Further details can be found in standard textbooks such as [70, 64].
@@ A.1 Forward and backward approximation
In numerical analysis, a “forward-error” often measures the distance between the true solution to a problem and the solution returned by a numerical algorithm. A related notion is the so-called “backward error” or “backward approximation.” For more details we can refer to the standard textbook of Higham [70]. In this case, the solution returned is the exact solution of a “nearby problem,” and the backward error quantifies the distance of the original problem to this nearby problem. In PCA, for example, a backward-approximation could be defined by $\widetilde{\boldsymbol{\mathrm{C}}}_{k}=\arg\min_{\boldsymbol{\mathrm{C}}^{\top}\boldsymbol{\mathrm{C}}=\boldsymbol{\mathrm{I}}_{k\times k}}\operatorname{tr}(\boldsymbol{\mathrm{C}}^{\top}\widetilde{\boldsymbol{\mathrm{H}}}\boldsymbol{\mathrm{C}})$. The backward error would be given $\|\boldsymbol{\mathrm{H}}-\widetilde{\boldsymbol{\mathrm{H}}}\|$ for some norm. An algorithm is backward stable if the backward error is always well-defined and bounded. A forward error type of approximation is often harder to achieve. A common rule-of-thumb states that
@@
However, this does not generally hold for eigenvalue problems. For example, Proposition 1.1 in [15] can be used to transformed a backward approximate diagonalization error to a forward error for the eigenvectors, but the bound depends on the minimum eigenvalue gap between any eigenvalue pair. As the authors point in Remark 1.2, special treatment is needed in terms of the invariant subspaces in the presence of multiple eigenvalues, which are thoroughly analyzed in this work.